Can someone help me on this problem?
I need t solve the system by addition:
x + y = 6
x - y = 4
Thank you,
Lisa

Answer by stanbon(75887)

x + y = 6
x - y = 4
---------------
Add the two quations to get:
2x=10
x=5
Substitute into the 1st equation to solve for "y":
5+y=6
y=1
SOLUTION:
x=5, y=1
Cheers,
Stan H.